Common salt obtained from seawater contains 8.775% NaCl by mass. The number of formula unit of NaCl present in 25 gram of the salt is:A. 3.367×10^23 formula unit B. 2.258×10^22 formula unit C. 3.176×10^23 formula unit D. 4.775×10^23 formula unit

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Akash
 
Let 100g of sea water contains 8.775g of sodium chloride
So 25g of water would have 2.19375g of sodium chloride
 
 
Molar mass of sodium chloride is 35.5+23=58.5g 
No of moles in 2.19375 g of sodium chloride is 2.19375/58.5 moles
No. Of molecules present would be (2.19375/58.5)*Avogadro number
 
 = 2.19375* 6* 10^23/(58.5)
 
 = 2.25 * 10^22 units
 
 
Regards
Arun (askIITians forum expert)
 
 
 
Regards
Arun (askIITians forum expert)